OpenStudy (anonymous):

(x - 4)2 = 36 Which of the following expressions represents the solutions to the given equation? -4 ± √(36) 4 ± √(36) 16 ± √(36)

OpenStudy (anonymous):

\[\huge (x - 4)^2 = 36\] \[\huge \rightarrow (x - 4) = \pm \sqrt{36}\] \[\huge \rightarrow x = 4\pm \sqrt{36}\] Hence \[\huge 4 \pm \sqrt{36}\] represents the solutions to the given equation. Thus option 2 is the correct answer.